# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

### v1

@@ 2.2. Orbifold Kähler compactification of $\mathbb{C}^{n}$
Our argument is similar to [Kol13, 3.1]. Let $\mathcal{C}^{\prime}$ denote the total space of the orbifold line bundle $\pi_{\mathcal{C}^{\prime}}:\mathcal{L}^{-1}\rightarrow\mathcal{Y}$. Then we have the identity $K_{\mathcal{C}^{\prime}}\otimes\mathcal{O}(\mathcal{Y}_{0})=\pi_{\mathcal{C}^{\prime}}^{*}K^{\mathrm{orb}}_{\mathcal{Y}}$. Away from the zero section $\mathcal{Y}_{0}$ we have $K_{\mathcal{C}^{\circ}}=\pi^{*}K^{\mathrm{orb}}_{\mathcal{Y}}=\pi^{*}(\mathcal{L}^{-r})\cong\mathcal{O}_{\mathcal{C}^{\circ}}$ where $\pi:\mathcal{C}^{\circ}\rightarrow\mathcal{Y}$ is the natural projection. Since $\mathcal{C}$ is normal, $K_{\mathcal{C}}$ is also trivial and hence $\mathcal{C}$ is Gorenstein.
@@
where $\mathcal{Y}_{0}$ is the zero section of $\mathcal{L}^{-1}$. Let $\mu_{2}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}^{\prime}$ be the resolution of singularities with simple normal crossing exceptional divisors $\{F_{i}\}$, we get:
@@
$=$ $\mu^{*}K_{\mathcal{C}}+r\widetilde{\mathcal{Y}}_{0}+\sum_{i}(a_{i}+r\cdot\mathrm{ord}_{F_{i}}(\mathcal{O}(-\mathcal{Y}_{0}))F_{i},$
where $\mathcal{O}(-\mathcal{Y}_{0})$ denotes the ideal sheaf of $\mathcal{Y}_{0}\subset\mathcal{C}^{\prime}$ and $\widetilde{\mathcal{Y}}_{0}$ is the strict transform of $\mathcal{Y}_{0}$ under $\mu_{2}$ and $\mu=\mu_{2}\circ\mu_{1}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ is a resolution of singularity. Note that $\mathcal{Y}_{0}$ has only quotient singularities which are klt. By inversion of adjunction (see [Kol04, Theorem 4.9] and the references there), $(\mathcal{C}^{\prime},\mathcal{Y}_{0})$ is plt which implies $a_{i}=a(F_{i},\mathcal{C}^{\prime},\mathcal{Y}_{0})>-1$. We see that $o\in\mathcal{C}$ is klt since the coefficients of the exceptional divisors of $\mu$ are bigger than $-1$. ∎
We can also write the minimal discrepancy of $o\in\mathcal{C}$ as
@@
We can then derive an algebraic formula for the minimal discrepancy of $o\in\mathcal{C}$ by adapting the argument as in [Kol13, Proof of Theorem 3.21] as follows.
First note that $\mathcal{C}^{\prime}$ has only quotient singularities that are contained in $\mathcal{Y}_{0}$. Let $p\in\mathcal{C}^{\prime}$ be a quotient singularity given by $\pi_{p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}/G_{p}$ where $G_{p}$ is a cyclic subgroup of $U(n)$. According to the local classification of smooth Seifert $\mathbb{C}^{*}$-bundles from [Kol04, 25], $\mathbb{C}^{n}/G_{p}$ is isomorphic to $\mathbb{C}^{n}/\frac{1}{m}(b_{1},\dots,b_{n})$ that satisfies the condition that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$ with $\{x_{1},\dots,x_{n}\}$ being linear coordinates on $\mathbb{C}^{n}$ and $b_{1}$ is relatively prime to $m$. We have the following formula for the minimal discrepancy of $o\in\mathcal{C}$ which can be seen as a generalization of Reid-Tai’s criterion for klt quotient singularities to the case of isolated Fano cone singularities.
@@
(14) $\mathrm{md}(o,\mathcal{C})=\min_{p,g}\left\{r,\frac{1}{m}\left(\sum_{i=1}^{n}w_{i}(g)+(r-1)w_{1}(g)\right)\right\}-1$
@@
We pick any exceptional divisor $F$ over $\mathbb{C}^{n}/G$ dominated by an exceptional divisor $E$ over $\mathbb{C}^{n}$ that is pointwise fixed by a cyclic subgroup $\langle g\rangle$ of order $k>1$. Assume $m=k\ell$. We have the equality $a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1=k^{-1}(a(E,\mathbb{C}^{n},\Delta^{\prime})+1)$ where $\Delta^{\prime}=(1-r)\pi_{p}^{*}\mathcal{Y}_{0}$. At the generic point of $E$ choose local coordinates $y_{1},\dots,y_{n}$ such that $E=\{y_{1}=0\}$, $g^{*}y_{1}=\epsilon^{\ell}y_{1}$ is a $g$-eigenfunction (up to a change of the generator of $\langle g\rangle$). Using the linear coordinates $\{x_{i}\}$ on $\mathbb{C}^{n}$ as above such that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$, we can write $x_{i}=y_{1}^{c_{i}}u_{i}$ where the $u_{i}$ are units. Then $g^{*}(dx_{1}\wedge\cdots\wedge dx_{n})$ vanishes to order $\geq(-1+\sum_{i}c_{i})$. If $g^{*}x_{i}=\epsilon^{w_{i}}x_{i}$ with $0\leq w_{i}<m$, then $\ell c_{i}\equiv w_{i}\text{ mod }m$. So we get:
$a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1$ $=$ $\frac{1}{k}(a(E,\mathbb{C}^{n},(1-r)\mathcal{Y}_{0})+1)$
$\geq$ $\frac{1}{k}(\sum_{i}c_{i}+(r-1)c_{1})\geq\frac{1}{k\ell}(\sum_{i}w_{i}+(r-1)w_{1}).$
Because of (13), this proves that the left-hand-side of (14) is greater than the right-hand-side.
Conversely any element $g\in G$ generates a cyclic subgroup which acts faithfully on $\mathbb{C}^{n}$ as $\mathbb{C}^{n}/\frac{1}{m}(w_{1},\dots,w_{n})$ where $\mathcal{Y}_{0}=\{x_{1}=0\}$. We want to show that $\frac{1}{m}(\sum_{i}w_{i}+(r-1)w_{1})\geq\mathrm{md}(o,\mathcal{C})$. If $\mathrm{gcd}(w_{1},\dots,w_{n})=\ell>1$, then the element $\frac{1}{m}(w_{1}/\ell,\cdots,w_{n}/\ell)$ contributes a smaller number to the right-hand-side. So we assume that $\mathrm{gcd}(w_{1},\dots,w_{n})=1$. Consider the weighted blow-up $\mathbb{C}^{n}$ with weights $(w_{1},\dots,w_{n})$. By [Kol13, pg. 104], a local chart is given by
$f:\mathbb{C}^{n}_{\mathbf{x}}/\frac{1}{w_{1}}(1,-w_{2},\dots,-w_{n})\rightarrow\mathbb{C}^{n}_{\mathbf{y}}$
@@
$a(F,\mathbb{C}^{n}/G,(1-r)\mathcal{Y}_{0})+1=\frac{a(\{z_{1}=0\},\mathbb{C}^{n},(1-r)\{z_{1}=0\})+1}{m}=\frac{1}{m}((r-1)w_{1}+\sum_{i=1}^{n}w_{i}).$
@@ 2.3. Conley-Zehnder indices and SFT degrees
We briefly review the basics of Conley-Zehnder indices and related concepts. Given a path $\phi:[0,1]\to\mathrm{Sp}(2n)$ in the symplectic matrices, such that $\phi(0)=\mathrm{id}$ and $\phi(1)\in\mathrm{Sp}^{*}(2n):=\left\{A\in\mathrm{Sp}(2n)\left|\det(A-\mathrm{id})\neq 0\right.\right\}$, we can define the integral Conley-Zehnder index $\mu_{\mathrm{CZ}}(\phi)$, which is characterized by the following properties [Gut14, Proposition 8]:
  1. (1)
@@
Properties (2), (4) and (6) determine Conley-Zehnder index uniqueness, see [Gut14, Proposition 9].
@@
  1. (a)
    Conley-Zehnder index has a generalization called the Robbin-Salamon index $\mu_{\mathrm{RS}}\in\frac{1}{2}\mathbb{Z}$ defined for any path in $\mathrm{Sp}(2n)$. One important feature of $\mu_{\mathrm{RS}}$ is that it is additive w.r.t. concatenation of paths [Gut14, Property 4 of Lemma 2 6]. By property (4) above, we have $\mu_{\mathrm{RS}}(\psi)=2\mu(\psi)$ for a loop $\psi$. Moreover, we have the following normalization formula for the path in $U(1)\subset Sp(2)$:
    (15) $\mu_{RS}(\{e^{i\theta};0\leq\theta\leq T\})=\left\{\begin{array}[]{ll}\frac{T}{\pi},&\text{if }T\in 2\pi\mathbb{Z}\\ 2\lfloor\frac{T}{2\pi}\rfloor+1,&\text{if }T\not\in 2\pi\mathbb{Z}.\end{array}\right.$
  2. (b)
@@
We now turn to the geometric setting. For a contact manifold $(M^{2n-1},\eta)$, let $\phi_{t}$ denote the flow generated by the Reeb vector field ${\mathfrak{v}}$. For a closed Reeb orbit $\gamma$, after choosing the trivialization of $\gamma^{*}(\xi:=\ker\eta)$, the linearized flow gives rise to a path in $\mathrm{Sp}(2n-2)$. By properties (2) and (4), the Conley-Zehnder index (including $\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) only depends on the induced trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, which will be denoted by $\mu_{\mathrm{CZ}}^{\tau}(\gamma)$ ($\mu^{\tau}_{\mathrm{LCZ}}(\gamma),\mu^{\tau}_{\mathrm{RS}}(\gamma)$) respectively. To get Conley-Zehnder indices for all Reeb orbits, we have the following:
@@
    For any Reeb orbit $\gamma$, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)\in{\mathbb{Z}_{2}}$ is well-defined. This is defined by $\mu_{\mathrm{LCZ}}^{\tau}(\gamma)\mod 2$ for any trivialization $\tau$ of $\det_{\mathbb{C}}\gamma^{*}\xi$. The well-definedness follows from Property (4).
@@
    If $c_{1}(\xi)=0$, $\det_{\mathbb{C}}\xi$ can be trivialized over $M$ (the set of trivializations is parameterized by $H^{1}(M;\mathbb{Z})$). After choosing a trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, we get $\mu^{\tau}_{\mathrm{LCZ}}(\gamma)$ for all Reeb orbits and it is independent of $\tau$ if $[\gamma]=0\in H_{1}(M;\mathbb{Q})$. In the case, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}(\gamma)\mod 2$.
@@
    If $c^{\mathbb{Q}}_{1}(\xi)=0$ in $H^{2}(M;\mathbb{Q})$, we can trivialize $\det_{\mathbb{C}}\oplus^{N}\xi$ for some $N\in\mathbb{Z}_{>0}$, after choosing a trivialization $\tau$, this allows us to define $\mu_{\mathrm{LCZ}}^{\mathbb{Q},\tau}(\gamma)=\mu_{\mathrm{LCZ}}^{\tau}(\oplus^{N}\phi)/N\in\mathbb{Q}$, where $\phi$ is the linearized flow along $\gamma$. This definition is independent of $N$ and is moreover independent of $\tau$ if $[\gamma]=0\in H_{1}(M;\mathbb{Q})$ [GZ21, §3.3.2]. This rational index coincides with the integer index above when $c_{1}(\xi)=0$. In general, one should think $(\mu^{\mathbb{Q}}_{\mathrm{LCZ}},\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}})\in\mathbb{Q}\times\mathbb{Z}_{2}$ is a pair of indices associated to an orbit.
@@
For a $12n-$ 1 dimensional contact manifold $(M.\xi)$ with rational first Chern class $c^{\mathbb{Q}}_{1}(\xi)=0$ and $H^{1}(M;\mathbb{Q})=0$, let $\gamma$ be a Reeb orbit for a contact form $\eta$, we define the lower SFT degree
$\mathrm{lSFT}(\gamma):=\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)+n-3.$
@@
$\mathrm{hmi}(M):=\sup_{\eta}\inf_{\gamma}\left\{\mathrm{lSFT}(\gamma)\right\}\in\mathbb{Q}\cup\{\pm\infty\},$
where $\eta$ is a contact form and $\gamma$ is a Reeb orbit of $\eta$. When $M$ is the contact link of an isolated $\mathbb{Q}$-Gorenstein normal singularity, McLean [McL16, Theorem 1.1] proved that $\mathrm{hmi}(M)$ is twice the minimal discrepancy number if $\mathrm{hmi}(M)\geq 0$ and $\mathrm{hmi}(M)<0$ is equivalent to that the minimal discrepancy number is negative.

### v3

@@ 2.2. Orbifold Kähler compactification of $\mathbb{C}^{n}$
The proof is similar to [Kol13, 3.1]. Let $\mathcal{C}^{\prime}$ denote the total space of the orbifold line bundle $\pi_{\mathcal{C}^{\prime}}:\mathcal{L}^{-1}\rightarrow\mathcal{Y}$. Then we have the identity $K_{\mathcal{C}^{\prime}}\otimes\mathcal{O}(\mathcal{Y}_{0})=\pi_{\mathcal{C}^{\prime}}^{*}K^{\mathrm{orb}}_{\mathcal{Y}}$. Away from the zero section $\mathcal{Y}_{0}$ we have $K_{\mathcal{C}^{\circ}}=\pi^{*}K^{\mathrm{orb}}_{\mathcal{Y}}=\pi^{*}(\mathcal{L}^{-r})\cong\mathcal{O}_{\mathcal{C}^{\circ}}$ where $\pi:\mathcal{C}^{\circ}\rightarrow\mathcal{Y}$ is the natural projection. Since $\mathcal{C}$ is normal, $K_{\mathcal{C}}$ is also trivial and hence $\mathcal{C}$ is Gorenstein.
@@
where $\mathcal{Y}_{0}$ is the zero section of $\mathcal{L}^{-1}$. Let $\mu_{2}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}^{\prime}$ be the resolution of singularities with simple normal crossing exceptional divisors $\{F_{i}\}$, we get an identity similar to (3):
@@
$=$ $\mu^{*}K_{\mathcal{C}}+r\widetilde{\mathcal{Y}}_{0}+\sum_{i}(a_{i}+r\cdot\mathrm{ord}_{F_{i}}(\mathcal{O}(-\mathcal{Y}_{0}))F_{i}.$
Here $\mathcal{O}(-\mathcal{Y}_{0})$ denotes the ideal sheaf of $\mathcal{Y}_{0}\subset\mathcal{C}^{\prime}$ and $\widetilde{\mathcal{Y}}_{0}$ is the strict transform of $\mathcal{Y}_{0}$ under $\mu_{2}$ and $\mu=\mu_{2}\circ\mu_{1}:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ is a resolution of singularity. The coefficient $a_{i}=a(F,\mathcal{C}^{\prime},\mathcal{Y}_{0})$ defined by the first equality is called the discrepancy of the exceptional divisor $F_{i}$ with respect to the pair $(\mathcal{C}^{\prime},\mathcal{Y}_{0})$ (see [Kol13, Definition 2.4]). Note that $\mathcal{Y}_{0}$ has only quotient singularities which are klt. By the theorem of inversion of adjunction (see [Kol13, Theorem 4.9]), $(\mathcal{C}^{\prime},\mathcal{Y}_{0})$ is plt (meaning “purely log terminal”, see [Kol13, Definition 2.8]) which implies $a_{i}=a(F_{i},\mathcal{C}^{\prime},\mathcal{Y}_{0})>-1$. So we get the formula:
$\mathrm{md}(o\in\mathcal{C})=\min\{r-1,a_{i}+r\cdot\mathrm{ord}_{F_{i}}(\mathcal{O}(-\mathcal{Y}_{0}))\}>-1$
which implies that $o\in\mathcal{C}$ is indeed klt (see Definition 2.2). ∎
By the linearity property of discrepancies (see [Kol13, Lemma 2.5]), $a(F_{i},\mathcal{C}^{\prime},(1-r)\mathcal{Y}_{0})=a(F_{i},\mathcal{C}^{\prime},\mathcal{Y}_{0})+r\cdot\mathrm{ord}_{F_{i}}(\mathcal{O}(-\mathcal{Y}_{0}))$. Therefore, we can also write the minimal discrepancy of $o\in\mathcal{C}$ as
@@
We can then derive an algebraic formula for the minimal discrepancy of $o\in\mathcal{C}$ by adapting the argument as in [Kol13, Proof of Theorem 3.21]. First note that $\mathcal{C}^{\prime}$ has only quotient singularities that are contained in $\mathcal{Y}_{0}$. Let $p\in\mathcal{C}^{\prime}$ be a quotient singularity given by $\pi_{p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}/G_{p}$ where $G_{p}$ is a cyclic subgroup of $U(n)$. According to the local classification of smooth Seifert $\mathbb{C}^{*}$-bundles from [Kol04, 25], $\mathbb{C}^{n}/G_{p}$ is isomorphic to $\mathbb{C}^{n}/\frac{1}{m}(b_{1},\dots,b_{n})$ that satisfies the condition that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$ with $\{x_{1},\dots,x_{n}\}$ being linear coordinates on $\mathbb{C}^{n}$ and $b_{1}$ is relatively prime to $m$. We have the following formula for the minimal discrepancy of $o\in\mathcal{C}$, which can be seen as a generalization of Reid-Tai’s criterion for klt quotient singularities to the case of isolated Fano cone singularities.
@@
(15) $\mathrm{md}(o,\mathcal{C})=\min_{p,g}\left\{r,\frac{1}{m}\left(rw_{1}(g)+\sum_{i=2}^{n}w_{i}(g)\right)\right\}-1$
@@
We pick any exceptional divisor $F$ over $\mathbb{C}^{n}/G_{p}$ dominated by an exceptional divisor $E$ over $\mathbb{C}^{n}$ that is pointwise fixed by a cyclic subgroup $\langle g\rangle$ of order $k>1$. Assume $m=k\ell$. We have the equality $a(F,\mathbb{C}^{n}/G_{p},(1-r)\mathcal{Y}_{0})+1=k^{-1}(a(E,\mathbb{C}^{n},\Delta^{\prime})+1)$ where $\Delta^{\prime}=(1-r)\pi_{p}^{*}\mathcal{Y}_{0}$ (see [Kol13, (2.42.4)]). At a general point of $E$ contained in a birational model over $X$, choose local coordinates $y_{1},\dots,y_{n}$ such that $E=\{y_{1}=0\}$, $g^{*}y_{1}=\epsilon^{\ell}y_{1}$ is a $g$-eigenfunction (up to a change of the generator of $\langle g\rangle$). Using the linear coordinates $\{x_{i}\}$ on $\mathbb{C}^{n}$ as above such that $\pi_{p}^{*}\mathcal{Y}_{0}=\{x_{1}=0\}$, we can write $g^{*}x_{i}=y_{1}^{c_{i}}u_{i}$ where the $u_{i}$ are units. Thus (the pullback of) $g^{*}(dx_{1}\wedge\cdots\wedge dx_{n})$ vanishes to order equal to $(-1+\sum_{i}c_{i})$ along $E$. By the definition of discrepancy (see remark 2.1), we get $a(E,\mathbb{C}^{n})=\sum_{i}c_{i}-1$, and by the linearity property $a(E,\mathbb{C}^{n},\Delta^{\prime})+1=a(E,\mathbb{C}^{n})+1+(r-1)c_{1}\geq\sum_{i}c_{i}+(r-1)c_{1}$. If $g^{*}x_{i}=\epsilon^{w_{i}}x_{i}$ with $0\leq w_{i}<m$, then $\ell c_{i}\equiv w_{i}\text{ mod }m$. So we get:
$a(F,\mathbb{C}^{n}/G_{p},(1-r)\mathcal{Y}_{0})+1$ $=$ $\frac{1}{k}(a(E,\mathbb{C}^{n},\Delta^{\prime})+1)$
$\geq$ $\frac{1}{k}(\sum_{i}c_{i}+(r-1)c_{1})\geq\frac{1}{k\ell}(\sum_{i=1}^{n}w_{i}+(r-1)w_{1})$
$=$ $\frac{1}{m}\left(rw_{1}(g)+\sum_{i=2}^{n}w_{i}(g)\right).$
Because of (14), this proves that the left-hand side of (15) is greater than the right-hand side.
Conversely any element $g\in G_{p}$ generates a cyclic subgroup which acts faithfully on $\mathbb{C}^{n}$ as $\mathbb{C}^{n}/\frac{1}{m}(w_{1},\dots,w_{n})$ where $\mathcal{Y}_{0}=\{x_{1}=0\}$. We want to show that $\frac{1}{m}(\sum_{i}w_{i}+(r-1)w_{1})\geq\mathrm{md}(o,\mathcal{C})+1$. If $\mathrm{gcd}(w_{1},\dots,w_{n})=\ell>1$, then the element $\frac{1}{m}(w_{1}/\ell,\cdots,w_{n}/\ell)$ contributes a smaller number to the right-hand-side. So we assume that $\mathrm{gcd}(w_{1},\dots,w_{n})=1$. Consider the weighted blow-up $\mathbb{C}^{n}$ with weights $(w_{1},\dots,w_{n})$. By [Kol13, pg. 106], a local chart is given by
$f:\mathbb{C}^{n}_{\mathbf{z}}/\frac{1}{w_{1}}(1,-w_{2},\dots,-w_{n})\rightarrow\mathbb{C}^{n}_{\mathbf{x}}$
@@
$a(F,\mathbb{C}^{n}/G_{p},(1-r)\mathcal{Y}_{0})+1=\frac{a(\{z_{1}=0\},\mathbb{C}^{n},(1-r)\{z_{1}=0\})+1}{m}=\frac{1}{m}((r-1)w_{1}+\sum_{i=1}^{n}w_{i}).$
@@
Example 2.17.
Set $(\mathcal{C},o)=(\mathbb{C}^{n},0)$ and let $\mathbb{C}^{*}$ act on $\mathbb{C}^{n}$ by $t\cdot(z_{1},\dots,z_{n})=(t^{a_{1}}z_{1},\dots,t^{a_{n}}z_{n})$ with $a_{i}\in\mathbb{Z}_{>0}$, $a_{1}\geq\dots\geq a_{n}$ and $\gcd(a_{1},\dots,a_{n})=1$. The corresponding orbifold $\mathcal{Y}=(\mathbb{C}^{n}\setminus 0)/\mathbb{C}^{*}=\mathbb{P}(a_{1},\dots,a_{n})$ is the weighted projective space with the Fano index $r=\sum_{j=1}^{n}a_{j}$ which is always greater than $n$ except for $\mathbb{P}^{n-1}=\mathbb{P}(1,\dots,1)$. It is easy to verify that the minimum on the right-hand-side of (15) is obtained when $p=[1,0,\dots,0]$ and $g=\epsilon=e^{2\pi\sqrt{-1}/a_{1}}\in G_{p}=\langle\epsilon\rangle\cong\mathbb{Z}_{a_{1}}$. Indeed, we have $w_{1}(g)=1$, $w_{i}(g)=a_{1}-a_{i}$ for $2\leq i\leq n$ and hence
$\frac{1}{m}\left(rw_{1}(g)+\sum_{i=2}^{n}w_{i}(g)\right)$ $=$ $\frac{1}{a_{1}}\left(\sum_{j=1}^{n}a_{j}+\sum_{i=2}^{n}(a_{1}-a_{i})\right)$
$=$ $n=\mathrm{md}(0,\mathbb{C}^{n})+1$
@@ 2.3. Conley-Zehnder indices and SFT degrees
[…]
  • (P1)
@@
Properties ((P2)), ((P4)) and ((P6)) determine Conley-Zehnder index uniquely, see [Gut14, Proposition 9].
@@
  • (P7)
    Conley-Zehnder index has a generalization called the Robbin-Salamon index $\mu_{\mathrm{RS}}\in\frac{1}{2}\mathbb{Z}$ defined for any path in $\mathrm{Sp}(2n)$. One important feature of $\mu_{\mathrm{RS}}$ is that it is additive w.r.t. concatenation of paths [Gut14, Property 4 of Lemma 26]. By property ((P4)) above, we have $\mu_{\mathrm{RS}}(\psi)=2\mu(\psi)$ for a loop $\psi$. Moreover, we have the following normalization formula for the path in $U(1)\subset Sp(2)$:
    (16) $\mu_{\mathrm{RS}}(\{e^{i\theta};0\leq\theta\leq T\})=\left\{\begin{array}[]{ll}\frac{T}{\pi},&\text{if }T\in 2\pi\mathbb{Z}\\ 2\lfloor\frac{T}{2\pi}\rfloor+1,&\text{if }T\not\in 2\pi\mathbb{Z}.\end{array}\right$.
  • (P8)
@@
We now turn to the geometric setting. For a contact manifold $(M^{2n-1},\eta)$, let $\phi_{t}$ denote the flow generated by the Reeb vector field $\mathfrak{v}$}. For a closed Reeb orbit $\gamma$, after choosing a trivialization of $\gamma^{*}(\xi:=\ker\eta)$, the linearized flow gives rise to a path in $\mathrm{Sp}(2n-2)$. By properties ((P2)) and ((P4)), the Conley-Zehnder index (including $\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) only depends on the induced trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, which will be denoted by $\mu_{\mathrm{CZ}}^{\tau}(\gamma)$ ($\mu^{\tau}_{\mathrm{LCZ}}(\gamma),\mu^{\tau}_{\mathrm{RS}}(\gamma)$) respectively. To get Conley-Zehnder indices for all Reeb orbits, we have the following:
@@
    For any Reeb orbit $\gamma$, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)\in{\mathbb{Z}_{2}}$ is well-defined. This is defined by $\mu_{\mathrm{LCZ}}^{\tau}(\gamma)\mod 2$ for any trivialization $\tau$ of $\det_{\mathbb{C}}\gamma^{*}\xi$. The well-definedness follows from Property ((P4)).
@@
    If $c_{1}(\xi)=0$, $\det_{\mathbb{C}}\xi$ can be trivialized over $M$ (the set of trivializations is parameterized by $H^{1}(M;\mathbb{Z})$). After choosing a trivialization $\tau$ of $\det_{\mathbb{C}}\xi$, we get $\mu^{\tau}_{\mathrm{LCZ}}(\gamma)$ for all Reeb orbits and it is independent of $\tau$ if $\gamma]=0\in H_{1}(M;\mathbb{Q})$. In the case, $\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}(\gamma)\mod 2$.
@@
    If $c^{\mathbb{Q}}_{1}(\xi)=0$ in $H^{2}(M;\mathbb{Q})$, we can trivialize $\det_{\mathbb{C}}\oplus^{N}\xi$ for some $N\in\mathbb{Z}_{>0}$, after choosing a trivialization $\tau$, this allows us to define $\mu_{\mathrm{LCZ}}^{\mathbb{Q},\tau}(\gamma)=\mu_{\mathrm{LCZ}}^{\tau}(\oplus^{N}\phi)/N\in\mathbb{Q}$, where $\phi$ is the linearized flow along $\gamma$. This definition is independent of $N$ and is moreover independent of $\tau$ if $\gamma]=0\in H_{1}(M;\mathbb{Q}) [$GZ21, §3.3.2]. This rational index coincides with the integer index above when $c_{1}(\xi)=0$. In general, one should think $(\mu^{\mathbb{Q}}_{\mathrm{LCZ}},\mu^{\mathbb{Z}_{2}}_{\mathrm{LCZ}})\in\mathbb{Q}\times\mathbb{Z}_{2}$ is a pair of indices associated to an orbit.
@@
Let $(M,\xi)$ be a 2n12n-1 dimensional contact manifold with rational first Chern class $c^{\mathbb{Q}}_{1}(\xi)=0$ and $H^{1}(M;\mathbb{Q})=0$. For any Reeb orbit $\gamma$ for a fixed contact form $\eta$, we define its lower SFT degree
$\mathrm{lSFT}_{\eta}(\gamma):=\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)+n-3.$
@@
(18) $\mathrm{hmi}(M):=\sup_{\eta}\inf_{\gamma}\left\{\mathrm{lSFT}_{\eta}(\gamma)\right\}\in\mathbb{R}\cup\{\pm\infty\},$
where $\eta$ ranges over all contact forms and $\gamma$ ranges over all Reeb orbits of $\eta$. When $M$ is the contact link of an isolated $\mathbb{Q}$-Gorenstein normal singularity, McLean [McL16, Theorem 1.1] proved that $\mathrm{hmi}(M)$ is twice the minimal discrepancy number if $\mathrm{hmi}(M)\geq 0$ and $\mathrm{hmi}(M)<0$ is equivalent to that the minimal discrepancy number is negative.